OpenStudy (anonymous):

The half-life of a bacteria is 250,000 years. What percent of a present amount will remain after 5000 years? So confused on how to set it up! The half-life of a bacteria is 250,000 years. What percent of a present amount will remain after 5000 years? So confused on how to set it up! @Mathematics

OpenStudy (anonymous):

start with \[A=(\frac{1}{2})^{\frac{t}{250,000}}\]

OpenStudy (anonymous):

then replace t by 5,000 to get \[(\frac{1}{2})^{\frac{5000}{250,000}}=(\frac{1}{2})^{.02}\]\]

OpenStudy (anonymous):

then a calculator. because you cannot do this by hand

OpenStudy (anonymous):

i get .986 rounded so about 98.6%

OpenStudy (anonymous):

ok awesome! That's what i was getting

OpenStudy (anonymous):

Thank you so much

OpenStudy (anonymous):

N/N(0)=e^-(ln2/250000)(5000)=0.9862x100=98.6% left...

OpenStudy (anonymous):

any question jsoeung?